Suppose there is economic contraction. (a) Show and explain how the economic performance will affect the prices of assets (bonds); (b) How will the price of assets impact the level of prospective de...

  1. Suppose there is economic contraction. (a) Show and explain how the economic performance will affect the prices of assets (bonds); (b) How will the price of assets impact the level of prospective deposits in your bank? Why? (c) Show and explain why the situation will be different during an economic expansion.

  2. (a) A customer of First National Bank (FNB) deposits $10,500 with FNB for 10 years. What is the annualized interest rate that the bank must pay if the customer’s deposit grows to $12,166.53 at the end of the ten-year term?

(b) Suppose the bank decides to pay 1.5% interest twice a year, what will be the future value amount that David Broke should receive?

3. Show and briefly explain the conditions under which an increase in non-borrowed reserves will

equilibrate the discount rate, short-term rate, and interest on reserves.
Answer #1

Initial investment is 10500$

number year = 10

returns with principal 12166.53$ on the 10th year

A) what is annualized interest rate paid?

total principal 10500 - to returns 12166.53 = 1666.53

1666.53/10 = 166.65 int amount per year

and int rate is 0.1587 for 10 years 1.587

B) if the bank decide to pay 1.5% twice in a six month

if its accumulated on only calculate on the initial principal

10500 * 1.5 = 157.5

= 157.5 * 20 (10 yrs)

= 3150

so the end of the 10th year, he will get 13650 $ with the int rate of 3.47%

3) option b would give him all benefits even he withdraw in short term since the bank pay interest in every six months that twice a year,